Calculate how much 20% alcohol solution and 60% alcohol solution must be mixed to end up with exactly 12 gallons of a 50% alcohol solution. You'll need ____ gallons of the 60% solution.
Answer by ankor@dixie-net.com(22740)

Calculate how much 20% alcohol solution and 60% alcohol solution must be mixed
to end up with exactly 12 gallons of a 50% alcohol solution.
You'll need ____ gallons of the 60% solution.
:
Let x = amt of 60% solution required
the total amt will be 12 gals, therefore:
(12-x) = amt of 20% solution
:
.60x + .20(12-x) = .50(12)
.60x + 2.4 - .20x = 6
.60x - .20x = 6 - 2.4
.40x = 3.6
x = 3.6/.4
x = 9 gal of 60% solution required
:
:
Check
.6(9) + .2(3) = 6